Find the derivative of the function. g(x) = 800(3 - e-0.1x)Find the derivative of the function. y = :etsFind the derivative of the function. y = 4e(x2 + 6)3Find the derivative of the function. y = In(e2x + 3)
Answer #1

= 4 ec?**)* (3)(2246) +(21) -0,1% 70) =800(3-6-0.18) 0/61) = 800 (0-(-0.1) 505 [01e01] 08 5 (0) 0% y=e 25 cas) (524) 5x4ees Y..+3) 29 € y a hace 2043 e 22 (e at han a (2-0), +0) 2a43 e 27 qe 2213 e